# Towards Self-Supervised FG-SBIR with Unified Sample Feature Alignment and Multi-Scale Token Recycling

### v3

@@ Abstract
Fine-Grained Sketch-Based Image Retrieval (FG-SBIR) aims to minimize the distance between sketches and corresponding images in the embedding space. However, scalability is hindered by the growing complexity of solutions, mainly due to the abstract nature of fine-grained sketches. In this paper, we propose an effective approach to narrow the gap between the two domains. It mainly facilitates unified mutual information sharing both intra- and inter-samples, rather than treating them as a single feature alignment problem between modalities. Specifically, our approach includes: (i) Employing dual weight-sharing networks to optimize alignment within the sketch and image domain, which also effectively mitigates model learning saturation issues. (ii) Introducing an objective optimization function based on contrastive loss to enhance the model’s ability to align features in both intra- and inter-samples. (iii) Presenting a self-supervised Multi-Scale Token Recycling (MSTR) Module featured by recycling discarded patch tokens in multi-scale features, further enhancing representation capability and retrieval performance. Our framework achieves excellent results on CNN- and ViT-based backbones. Extensive experiments demonstrate its superiority over existing methods. We also introduce Cloths-V1, the first professional fashion sketch-image dataset, utilized to validate our method and will be beneficial for other applications.
@@ Introduction
In this study, we introduce the ARNet (Feature Alignment and Token Recycling). We experimentally discovered that a straightforward methodology can achieve favorable results and enhance the model’s versatility by switching any backbone. Additionally, recycling token features can improve the model’s feature representation capability. Specifically, we simultaneously optimize feature distribution alignment intra- and inter-samples through structural design and loss function configuration. Our initial experiments revealed that single-branch encoders can tend to train saturation. Therefore, we introduced a dual weight-sharing network structure to enhance feature consistency within the sketch and image domains, ensuring effective learning of key fine-grained features. We also proposed a new objective optimization method for self-supervised training, promoting optimization of features both intra- and inter-samples, as shown in Fig. 1 (c), which differs from traditional Triplet Loss (a) and Contrastive Loss (b). Furthermore, we introduced the plug-and-play Multi-Scale Token Recycling (MSTR) Module to recycle discarded Patch Tokens and used Multi-Scale Token-Wise Contrastive Learning to filter out similar features between tokens while retaining unique features. This approach enhances feature learning from individual samples, thereby improving the model’s representation ability.
@@
Contrastive learning improves model representational ability by maximizing positive sample similarity and minimizing negative sample similarity. In earlier works, (Wu et al. 2018) proposed the memory bank to store data features. Since then, methods like (Ye et al. 2019; Tian, Krishnan, and Isola 2020; Chen et al. 2020) used data augmentation, learned from multiple views as input or larger pos-neg sample pairs, etc. (He et al. 2020) proposed the momentum encoder to address sample selection bias. (Caron et al. 2020) used a clustering algorithm to reduce computational costs. Recently,  (Wang et al. 2021b; Tang et al. 2020, 2023) proposed pixel contrast for image segmentation or synthesis. These methods drive unsupervised representation learning to achieve superior performance. Similarly, we can leverage the concept of contrastive learning to FG-SBIR for better retrieval performance.
@@
Vision Transformer (ViT) (Dosovitskiy et al. 2020) uses a self-attention mechanism like (Vaswani et al. 2017), leading to a more effective way to model global features compared to traditional CNNs. Based on this, methods such as (Han et al. 2021; Liu et al. 2021; Wang et al. 2021a) proposed smaller patches, the shift-window mechanism, progressive shrinking pyramid, etc. to further enhance the capability of ViTs. However, only a few studies (Abhra et al. 2022; Sain et al. 2023) have explored ViT in the field of FG-SBIR. Moreover, these methods choose the [CLS] Token as feature output and discard other Patch Tokens, which will also result in discarding some potentially useful features. Therefore, we introduce the ViT into FG-SBIR and further improve model performance by recycling the discarded features.
@@
We begin with an initial structure as shown in Fig. 2 (a), using the prevalent pre-trained ViT-B/16 (Dosovitskiy et al. 2020) as our backbone and modifying the dimensions of the classification layer. However, we encountered the first challenge: the model’s performance exhibited saturation around the 150-epoch mark, and subsequent training iterations yielded negligible enhancements, as illustrated by the A0 curve in Fig. 2 (b)  [^1] A0: no augmentation. A1: image scaling, rotation, and cropping operations. A2: similar to Moco (He et al. 2020).. We tried experimenting with conventional solutions like data augmentation, output feature normalization, etc. Unfortunately, these methods did not show improvements. In terms of data augmentation, methods like A1 and A2 appeared to have no benefit in the generalization capability or even drive the model into overfitting territory, as shown in Fig. 2 (b) by the A1 and A2 curves.
@@
We begin by randomly loading sketch-image sample pairs from the dataset and performing data augmentation. The augmented and original samples are then input into the weight-sharing network $\mathcal{E}$ as previously described. Finally, we utilize a linear layer to map the [CLS] Token of sketch and image features to a joint embedding space. Specifically, denoting the sketches and images within our dataset as $\mathcal{S}$ and $\mathcal{I}$ respectively, with the number of samples in each minibatch designated as m, we succinctly represent our data-loading process as $\mathcal{B}_{m}\Rightarrow\{s_{j}\in\mathcal{S},i_{j}\in\mathcal{I}\mid s_{j}\text{ and }i_{j}\in\mathcal{R}^{224\times 224\times 3}\}$, and $j=1,2,\dots,m$. Initiating the feature extraction process, we begin by applying data augmentation to the input samples $\mathcal{B}_{m}$, generating the sample pairs denoted as $\mathcal{B}^{\prime}_{m}\Rightarrow\{(s_{j},s^{\prime}_{j},i_{j},i^{\prime}_{j})\mid s_{j}\in\mathcal{S},i_{j}\in\mathcal{I},s^{\prime}_{j}\in\mathcal{S}\text{ and }i^{\prime}_{j}\in\mathcal{I}\text{are augmented samples}\}$. Subsequently, the augmented samples undergo patch encoding and positional encoding before being fed to the encoders $\mathcal{E}_{skt}$ and $\mathcal{E}_{img}$, obtaining the sketch features $\mathcal{F}_{skt}\Rightarrow\{f_{skt}\leftarrow\mathcal{E}_{skt}(s),f^{\prime}_{skt}\leftarrow\mathcal{E}_{skt}(s^{\prime})\}$ and image features $\mathcal{F}_{img}\Rightarrow\{f_{img}\leftarrow\mathcal{E}_{img}(i),f^{\prime}_{img}\leftarrow\mathcal{E}_{img}(i^{\prime})\}$ for the input samples $\mathcal{B}_{m}$. Finally, the resulting features $\mathcal{F}_{skt}$ and $\mathcal{F}_{img}$ are channeled through a linear layer, mapping them into the $\mathcal{R}^{1\times 512}$ dimension within a joint embedding space. This process yields the embedded sketch features $\mathcal{F}_{skt}^{out}$ and the image features $\mathcal{F}_{img}^{out}$ for retrieval in a straightforward way.
@@
Specifically, Clothes-V1 has 1200 (500) sketches (images), containing 925 (380) and 275 (120) for the training and validation set. Each image corresponds to at least one sketch and at most five sketches. The dataset is further categorized into types such as sweaters, shirts, jackets, dresses, cheongsams, etc. Moreover, we are actively working to continuously expand our dataset as soon as possible.
@@
We compare our approach with 12 state-of-the-art methods: Triplet-SN (Yu et al. 2016) adopts triplet loss for model training. Triplet-Att-SN (Song et al. 2017) introduces an attention module to improve detailed features. OnTheFly (Bhunia et al. 2020) employs reinforcement learning and multi-stage sketch input to retrieve more quickly. B-Siamese (Sain et al. 2020) is builds on Triplet-SN, but with a stronger backbone. CMHM-SBIR (Sain et al. 2020) proposes a co-attention module to model cross-modal multi-level structures. SketchAA (Yang et al. 2021) proposes a multi-granularity modeling approach to represent sketch appearance and structural features. Semi-Sup (Bhunia et al. 2021) introduces a semi-supervised framework for cross-modal retrieval. StyleMeUp (Sain et al. 2021) proposes a meta-learning framework for generalizing to unseen sketch styles. Adpt-SBIR (Bhunia et al. 2022b) introduces a model-agnostic meta-learning framework. Part-SBIR (Chowdhury et al. 2022) proposes cross-modal domain associations in a part-aware manner using optimal transport. NT-SBIR (Bhunia et al. 2022a) proposes a framework for sketch stroke subset selection based on reinforcement learning. EUPS-SBIR (Sain et al. 2023) proposes using a triplet loss within a modality and pre-training a teacher model on a large set of unlabeled photos.
@@
As presented in Tables 1 and 2, our proposed method exhibits superior performance compared to other baselines on three datasets. (i) $\mathtt{QMUL\text{-}Chair\text{-}V2}$: Our model achieves new SOTA results and its Top5 accuracy can surpass EUPS-SBIR’s Top10 accuracy. For Top1 accuracy, it outperforms nearly 10% compared to NT-SBIR and Part-SBIR. (ii) $\mathtt{QMUL\text{-}Shoe\text{-}V2}$: Although our model does not secure the top ranking, it still demonstrates encouraging potential with third position, boasting a slight gap in three metrics compared to the existing methods. This can be attributed to the dataset’s larger size and the relatively lower specificity of shoe sketches compared to the chair and clothes, leading to a more closely clustered within the joint embedding space. (iii) $\mathtt{Clothes\text{-}V1}$: The results present a clear advantage of our model in Top1 and Top5 accuracy, along with an impressive nearly 100% accuracy in Top10. (iv) $\mathtt{ARNet~{}Full}$: Based on the ARNet Basic, our model achieved excellent results on three datasets with steady improvement by our proposed MSTR Module. These experimental results demonstrate our model’s remarkable retrieval capabilities in a straightforward way, as well as the effectiveness of our approach for recycling discarded Patch Tokens.
@@
We compare three different data processing strategies using two data loading methods based on the Contrast loss $\mathtt{T}$ and the Triplet loss $\mathtt{T}$ in Table 3. The results demonstrated that the $\mathtt{C}$-based method consistently outperforms the $\mathtt{T}$-based method under the same conditions. This superiority is attributed to our loss function’s comprehensive in-sample and inter-sample optimization, which provides greater gradient contrast information in each mini-batch optimization, thereby enhancing overall performance.
Table 4: Ablation Studies on MSTR Module on QMUL-Chair-V2. ($\mathtt{max}$, $\mathtt{mean}$, and $\mathtt{cat}$: Use MaxPool, AvgPool, and Concatenation for Discarded Patch Tokens.)
ID Recycle Map Layer Dim QMUL-Chair-V2(%)
@@
As shown in Table 4, we conducted ablation experiments on the MSTR module from four perspectives: recycling discarded Patch Tokens (Recycle), constructing $\mathtt{MAP}$ (Map), the number of Transformer Layers (Layer), and the feature dimension used for Map (Dim). The results indicate that simple methods like MaxPool, AvgPool, and Concatenation not only fail to improve but may even negatively impact model performance. Using the map can enhance contrastive learning between tokens, thus improving model performance. When examining Layer, 2 layers outperformed both 1 and 3 layers; 1 layer fails to encode feature information adequately, while 3 layers result in over-encoding and negative effects. Regarding Dim, we observed the same phenomenon that 8 cannot encapsulate all unique information, while 24 and 48 include too much irrelevant information, weakening the feature representation of discarded Patch Tokens. These findings collectively demonstrate the effectiveness of our MSTR module.

### v4

# ARNet: Self-Supervised FG-SBIR with Unified Sample Feature Alignment and Multi-Scale Token Recycling

@@ Abstract
Fine-Grained Sketch-Based Image Retrieval (FG-SBIR) aims to minimize the distance between sketches and corresponding images in the embedding space. However, scalability is hindered by the growing complexity of solutions, mainly due to the abstract nature of fine-grained sketches. In this paper, we propose an effective approach to narrow the gap between the two domains. It mainly facilitates unified mutual information sharing both intra- and inter-samples, rather than treating them as a single feature alignment problem between modalities. Specifically, our approach includes: (i) Employing dual weight-sharing networks to optimize alignment within the sketch and image domain, which also effectively mitigates model learning saturation issues. (ii) Introducing an objective optimization function based on contrastive loss to enhance the model’s ability to align features in both intra- and inter-samples. (iii) Presenting a self-supervised Multi-Scale Token Recycling (MSTR) Module featured by recycling discarded patch tokens in multi-scale features, further enhancing representation capability and retrieval performance. Our framework achieves excellent results on CNN- and ViT-based backbones. Extensive experiments demonstrate its superiority over existing methods. We also introduce Cloths-V1, the first professional fashion sketch-image dataset, utilized to validate our method and will be beneficial for other applications. Our code, new datasets, and pre-trained model are available at https://github.com/ExponentiAI/ARNet.
11footnotetext: Corresponding Authors.
@@ Introduction
In this study, we introduce the ARNet (Feature Alignment and Token Recycling). We experimentally discovered that a straightforward methodology can achieve favorable results and enhance the model’s versatility by switching any backbone. Additionally, recycling token features can improve the model’s feature representation capability. Specifically, we simultaneously optimize feature distribution alignment intra- and inter-samples through structural design and loss function configuration. Our initial experiments revealed that single-branch encoders can tend to train saturation. Therefore, we introduced a dual weight-sharing network structure to enhance feature consistency within the sketch and image domains, ensuring effective learning of key fine-grained features. We also proposed a new objective optimization method for self-supervised training, promoting optimization of features in both intra- and inter-samples, as shown in Fig. 1 (c), which differs from traditional Triplet Loss (a) and Contrastive Loss (b). Furthermore, we introduced the plug-and-play Multi-Scale Token Recycling (MSTR) Module to recycle discarded Patch Tokens and used Multi-Scale Token-Wise Contrastive Learning to filter out similar features between tokens while retaining unique features. This approach enhances feature learning from individual samples, thereby improving the model’s representation ability.
@@
Contrastive learning improves model representational ability by maximizing positive sample similarity and minimizing negative sample similarity. In earlier works, Wu et al. (2018) proposed the memory bank to store data features. Since then, methods like (Ye et al. 2019; Tian, Krishnan, and Isola 2020; Chen et al. 2020; duan2024mining) used data augmentation, learned from multiple views as input or larger pos-neg sample pairs, etc. He et al. (2020) proposed the momentum encoder to address sample selection bias. Caron et al. (2020) used a clustering algorithm to reduce computational costs. Recently, Wang et al. (2021b); Tang et al. (2020, 2023) proposed pixel contrast for image segmentation or synthesis. These methods drive unsupervised representation learning to achieve superior performance. Similarly, we can leverage the concept of contrastive learning to FG-SBIR for better retrieval performance.
@@
Vision Transformer (ViT) (Dosovitskiy et al. 2020) uses a self-attention mechanism like (Vaswani et al. 2017), leading to a more effective way to model global features compared to traditional CNNs. Based on this, methods such as (Han et al. 2021; thiry2024towards; Liu et al. 2021; Wang et al. 2021a) proposed smaller patches, the shift-window mechanism, progressive shrinking pyramid, etc. to further enhance the capability of ViTs. However, only a few studies (Abhra et al. 2022; Sain et al. 2023) have explored ViT in the field of FG-SBIR. Moreover, these methods choose the [CLS] Token as feature output and discard other Patch Tokens, which will also result in discarding some potentially useful features. Therefore, we introduce the ViT into FG-SBIR and further improve model performance by recycling the discarded features.
@@
We begin with an initial structure as shown in Fig. 2 (a), using the prevalent pre-trained ViT-B/16 (Dosovitskiy et al. 2020) as our backbone and modifying the dimensions of the classification layer. However, we encountered the first challenge: the model’s performance exhibited saturation around the 150-epoch mark, and subsequent training iterations yielded negligible enhancements, as illustrated by the A0 curve in Fig. 2 (b) ***A0: no augmentation. A1: image scaling, rotation, and cropping operations. A2: similar to Moco (He et al. 2020).. We tried experimenting with conventional solutions like data augmentation, output feature normalization, etc. Unfortunately, these methods did not show improvements. In terms of data augmentation, methods like A1 and A2 appeared to have no benefit in the generalization capability or even drive the model into overfitting territory, as shown in Fig. 2 (b) by the A1 and A2 curves.
@@
We begin by randomly loading sketch-image sample pairs from the dataset and performing data augmentation. The augmented and original samples are then input into the weight-sharing network $\mathcal{E}$ as previously described. Finally, we utilize a linear layer to map the [CLS] Token of sketch and image features to a joint embedding space. Specifically, denoting the sketches and images within our dataset as $\mathcal{S}$ and $\mathcal{I}$ respectively, with the number of samples in each minibatch designated as m, we succinctly represent our data-loading process as $\mathcal{B}_{m}\Rightarrow\{s_{j}\in\mathcal{S},i_{j}\in\mathcal{I}\mid s_{j}\text{ and }i_{j}\in\mathcal{R}^{224\times 224\times 3}\}$, and $j=1,2,\dots,m$. Initiating the feature extraction process, we begin by applying data augmentation to the input samples $\mathcal{B}_{m}$, generating the sample pairs denoted as $\mathcal{B}^{\prime}_{m}\Rightarrow\{(s_{j},s^{\prime}_{j},i_{j},i^{\prime}_{j})\mid s_{j}\in\mathcal{S},i_{j}\in\mathcal{I},s^{\prime}_{j}\in\mathcal{S}\text{ and }i^{\prime}_{j}\in\mathcal{I}\text{ are augmented samples}\}$. Subsequently, the augmented samples undergo patch encoding and positional encoding before being fed to the encoders $\mathcal{E}_{skt}$ and $\mathcal{E}_{img}$, obtaining the sketch features $\mathcal{F}_{skt}\Rightarrow\{f_{skt}\leftarrow\mathcal{E}_{skt}(s),f^{\prime}_{skt}\leftarrow\mathcal{E}_{skt}(s^{\prime})\}$ and image features $\mathcal{F}_{img}\Rightarrow\{f_{img}\leftarrow\mathcal{E}_{img}(i),f^{\prime}_{img}\leftarrow\mathcal{E}_{img}(i^{\prime})\}$ for the input samples $\mathcal{B}_{m}$. Finally, the resulting features $\mathcal{F}_{skt}$ and $\mathcal{F}_{img}$ are channeled through a linear layer, mapping them into the $\mathcal{R}^{1\times 512}$ dimension within a joint embedding space. This process yields the embedded sketch features $\mathcal{F}_{skt}^{out}$ and image features $\mathcal{F}_{img}^{out}$ for retrieval in a straightforward way.
@@
Specifically, Clothes-V1 has 1200 (500) sketches (images), containing 925 (380) and 275 (120) for the training and validation set. Each image corresponds to at least one sketch and at most five sketches. The dataset is further categorized into types such as sweaters, shirts, jackets, dresses, cheongsams, etc. Please refer to our GitHub for more details.
@@
We compare our approach with 12 state-of-the-art methods: Triplet-SN (Yu et al. 2016) adopts triplet loss for model training. Triplet-Att-SN (Song et al. 2017) introduces an attention module. OnTheFly (Bhunia et al. 2020) employs reinforcement learning and multi-stage sketch input. B-Siamese (Sain et al. 2020) is builds on Triplet-SN, but with a stronger backbone. CMHM-SBIR (Sain et al. 2020) proposes a cross-modal multi-level structure. SketchAA (Yang et al. 2021) proposes a multi-granularity modeling approach. Semi-Sup (Bhunia et al. 2021) introduces a semi-supervised framework for cross-modal retrieval. StyleMeUp (Sain et al. 2021) proposes a meta-learning framework for unseen sketch styles. Adpt-SBIR (Bhunia et al. 2022b) introduces a model-agnostic meta-learning framework. Part-SBIR (Chowdhury et al. 2022) proposes cross-modal domain associations by optimal transporting. NT-SBIR (Bhunia et al. 2022a) uses reinforcement learning for sketch stroke subset selection. EUPS-SBIR (Sain et al. 2023) employs triplet loss and pre-trained teacher model.
@@
As presented in Tables 1 and 2, our proposed method exhibits superior performance compared to other baselines on three datasets. (i) $\mathtt{QMUL\text{-}Chair\text{-}V2}$: Our model achieves new SOTA results and its Top5 accuracy can surpass EUPS-SBIR’s Top10 accuracy. For Top1 accuracy, it outperforms nearly 10% compared to NT-SBIR and Part-SBIR. (ii) $\mathtt{QMUL\text{-}Shoe\text{-}V2}$: Although our model does not secure the top ranking, it still demonstrates encouraging potential with third position, boasting a slight gap in three metrics compared to the existing methods. This can be attributed to the dataset’s larger size and the more relatively irregular sketch strokes compared to the chair and clothes, leading to more complex mappings and being closely clustered within the joint embedding space, which may affect the optimization process. (iii) $\mathtt{Clothes\text{-}V1}$: The results present a clear advantage of our model in Top1 and Top5 accuracy, along with an impressive nearly 100% accuracy in Top10. (iv) $\mathtt{ARNet~{}Full}$: Based on the ARNet Basic, our model achieved excellent results on three datasets with steady improvement by our proposed MSTR Module. These experimental results demonstrate our model’s remarkable retrieval capabilities in a straightforward way, as well as the effectiveness of our approach for recycling discarded Patch Tokens.
@@
We compare three different data processing strategies using two data loading methods based on the Contrast loss $\mathtt{C}$ and the Triplet loss $\mathtt{T}$ in Table 3. The results demonstrated that the $\mathtt{C}$-based method consistently outperforms the $\mathtt{T}$-based method under the same conditions. This superiority is attributed to our loss function’s comprehensive in-sample and inter-sample optimization, which provides greater gradient contrast information in each mini-batch optimization, thereby enhancing overall performance.
Table 4: Ablation Studies on our MSTR Module on QMUL-Chair-V2. ($\mathtt{max}$, $\mathtt{mean}$, and $\mathtt{cat}$: Use MaxPool, AvgPool, and Concatenation for Discarded Patch Tokens.)
ID Rcyc Map LayT LayS Dim QMUL-Chair-V2(%)
@@
As shown in Table 4, we conducted ablation experiments on the MSTR module from four perspectives: recycling discarded Patch Tokens (Rcyc), constructing $\mathtt{MAP}$ (Map), the number of Transformer Layers (LayT), the number of feature Scale Layers (LayS), and the feature dimension used for Map (Dim). The results indicate that simple methods like MaxPool, AvgPool, and Concatenation not only fail to improve but may even negatively impact model performance. Using the map can enhance contrastive learning between tokens, thus improving model performance. When examining LayT, 2 layers outperformed both 1 and 3 layers. This is mainly because 1 layer fails to encode feature information adequately, while 3 layers result in over-encoding and negative effects. For validating LayS, we evaluated models with 2 and 4 layers compared to our 3-layer setting. The results demonstrate that using 3 layers improves Top1 accuracy by approximately 2% and 3% compared to the settings of 2 and 4 layers, while also achieving balanced performance in Top5 and Top10 accuracy. Regarding Dim, we observed a similar phenomenon that 8 cannot encapsulate all unique information, while 24 and 48 include too much irrelevant information, weakening the feature representation of discarded Patch Tokens. These findings collectively demonstrate the effectiveness of our MSTR module.